# Continuous Variable Structured Collision Models

### v1

@@ 4 Conclusions And Outlook
This work presents a general method of designing a collision model that includes structure within the environmental units using continuous-variable quantum systems, namely harmonic oscillators. In some cases, it is possible to construct a Lyapunov equation for the continuous-time evolution of the system’s covariance matrix. We demonstrated the effect of an internal structure within the environment may have over the system’s dynamics and thermodynamics, particularly with respect to the effective temperature of the environmental units as a function of their internal coupling strength $\lambda_{I}$. The dynamics of a system interacting through beamsplitter-like interactions via a collision model with units composed of spring-like coupled oscillators, initially prepared in a thermal state, were found to be analogous to the dynamics due to interactions with single oscillator environmental units prepared in a squeezed thermal state. The magnitude of this squeezing also corresponded to a function of the internal coupling strength. This effective squeezing arises naturally from environmental units prepared in thermal equilibrium, thus considered free within the resource theory of quantum thermodynamics. This is in stark contrast to previous works which considered environmental units which are artificially prepared in squeezed states, thus necessarily employing additional thermodynamic resources.
@@
Verification of these results may be currently achievable through various experimental methods, such as through trapped ions or photonic systems. The effect of the topology of the structure within the environmental units would be of keen interest for further study, alongside the effects of varying the initial states, interactions between the system and environment, and of course the internal interactions within the environmental units. Finally, it would be of practical interest to examine the effects of a structured environment to the performances of heat engines and refrigerators, as well as in the charging process of quantum batteries.
@@ Appendix A Superoperator Derivation
The derivation of the superoperators relating the total dynamics of the system and environment to the reduced dynamics of solely the system, $\mathcal{L}_{0}(\sigma_{\text{tot}}(t))$, $\mathcal{L}_{1}(\sigma_{\text{tot}}(t))$, $\mathcal{L}_{2}(\sigma_{\text{tot}}(t))$ can be found by considering the set of elements $i$, $j$ of the matrices produced by the total dynamic operators $\mathcal{U}_{0}(\sigma_{\text{tot}}(t))$, $\mathcal{U}_{1}(\sigma_{\text{tot}}(t))$, $\mathcal{U}_{2}(\sigma_{\text{tot}}(t))$, for which $i=1,2...,2N_{S}$ and $j=1,2...,2N_{S}$. More intuitively, one can consider this as selecting only the “top-left” $2N_{S}\crossproduct 2N_{S}$ block of the total system-environment covariance matrix after applying the total evolution operators. The action of extracting the properties and dynamics of solely the system will be denoted here as $[\cdot]_{S}$, such that $\mathcal{L}_{i}(\sigma_{\text{tot}}(t))=[\mathcal{U}_{i}(\sigma_{\text{tot}}(t))]_{S}$.
@@
One can treat the $D_{SE}$ terms using the notation of $[\cdot]_{S}$, retaining only the $[D_{SE}\sigma_{\text{tot}}(t)]_{S}$ and $[\sigma_{\text{tot}}(t)D_{SE}^{\intercal}]_{S}$ contributions. Similarly, applying this operation on the $D_{S}^{\prime}$ terms produces $[D^{\prime}_{S}\sigma_{S}(t)]_{S}=D_{S}\sigma_{S}(t)$ and $[\sigma_{\text{tot}}(t)D_{S}^{\prime\intercal}]_{S}=\sigma_{S}(t)D_{S}^{\intercal}$. The dynamics of the system due to the first order $\delta t$ terms can then be written as
@@
$\sum_{k}D_{SE_{ik}}\sigma_{\text{tot}_{kl}}(t){D^{\prime}}^{\intercal}_{E_{lj}}=0;\qquad 1\leq j\leq 2N_{S}.$ (103)
@@ Appendix B Environmental Unit Normal-Mode Decomposition
$H_{E_{x}}=\frac{1}{N_{E}}\sum_{j,m,m^{\prime}}e^{-ij(k_{m}+k_{m^{\prime}})}\left[\frac{\omega_{E}^{2}}{2}\tilde{x}_{m}\tilde{x}_{m^{\prime}}+\lambda_{I}(1+e^{i(k_{m}+k_{m^{\prime}})}-e^{-ik_{m}}-e^{-ik_{m^{\prime}}})\tilde{x}_{m}\tilde{x}_{m^{\prime}}\right].$ (110)

### v2

@@ 4 Conclusions And Outlook
This work presents a general method of designing a collision model that includes structure within the environmental units using continuous-variable quantum systems, namely harmonic oscillators. In some cases, it is possible to construct a Lyapunov equation for the continuous-time evolution of the system’s covariance matrix. We demonstrated the effect of an internal structure within the environment may have over the system’s dynamics and thermodynamics, particularly with respect to the effective temperature of the environmental units as a function of their internal coupling strength $\lambda_{I}$. The dynamics of a system interacting through beamsplitter-like interactions via a collision model with units composed of spring-like coupled oscillators, initially prepared in a thermal state, were found to be analogous to the dynamics due to interactions with single oscillator environmental units prepared in a squeezed thermal state. The magnitude of this squeezing also corresponded to a function of the internal coupling strength. This effective squeezing arises naturally from environmental units prepared in thermal equilibrium, thus considered free within the resource theory of quantum thermodynamics. This is in stark contrast to previous works which considered environmental units which are artificially prepared in squeezed states [32, 53, 54], thus necessarily employing additional thermodynamic resources.
@@
Verification of these results may be currently achievable through various experimental methods, such as through trapped ions [55, 56, 57] or photonic systems [58, 59]. The effect of the topology of the structure within the environmental units would be of keen interest for further study, alongside the effects of varying the initial states, interactions between the system and environment, and of course the internal interactions within the environmental units. Finally, it would be of practical interest to examine the effects of a structured environment to the performances of heat engines and refrigerators, as well as in the charging process of quantum batteries.
@@ Appendix A Superoperator Derivation
The derivation of the superoperators relating the total dynamics of the system and environment to the reduced dynamics of solely the system, $\mathcal{L}_{0}(\sigma_{\text{tot}}(t))$, $\mathcal{L}_{1}(\sigma_{\text{tot}}(t))$, $\mathcal{L}_{2}(\sigma_{\text{tot}}(t))$ can be found by considering the set of elements $i$, $j$ of the matrices produced by the total dynamic operators $\mathcal{U}_{0}(\sigma_{\text{tot}}(t))$, $\mathcal{U}_{1}(\sigma_{\text{tot}}(t))$, $\mathcal{U}_{2}(\sigma_{\text{tot}}(t))$, for which $i=1,2...,2N_{S}$ and $j=1,2...,2N_{S}$. More intuitively, one can consider this as selecting only the “top-left” $2N_{S}\times 2N_{S}$ block of the total system-environment covariance matrix after applying the total evolution operators. The action of extracting the properties and dynamics of solely the system will be denoted here as $\cdot]_{S}$, such that $\mathcal{L}_{i}(\sigma_{\text{tot}}(t))=[\mathcal{U}_{i}(\sigma_{\text{tot}}(t))]_{S}$.
@@
One can treat the $D_{SE}$ terms using the notation of $\cdot]_{S}$, retaining only the $D_{SE}\sigma_{\text{tot}}(t)]_{S}$ and $\sigma_{\text{tot}}(t)D_{SE}^{\intercal}]_{S}$ contributions. Similarly, applying this operation on the $D_{S}^{\prime}$ terms produces $D^{\prime}_{S}\sigma_{S}(t)]_{S}=D_{S}\sigma_{S}(t)$ and $\sigma_{\text{tot}}(t)D_{S}^{\prime\intercal}]_{S}=\sigma_{S}(t)D_{S}^{\intercal}$. The dynamics of the system due to the first order $\delta t$ terms can then be written as
@@
$\sum_{kl}D_{SE_{ik}}\sigma_{\text{tot}_{kl}}(t){D^{\prime}}^{\intercal}_{E_{lj}}=0;\qquad 1\leq j\leq 2N_{S}.$ (103)
@@ Appendix B Environmental Unit Normal-Mode Decomposition
$H_{E_{x}}=\frac{1}{N_{E}}\sum_{j,m,m^{\prime}}e^{-ij(k_{m}+k_{m^{\prime}})}\left[\frac{\omega_{E}^{2}}{2}\tilde{x}_{m}\tilde{x}_{m^{\prime}}+\lambda_{I}(1+e^{-i(k_{m}+k_{m^{\prime}})}-e^{-ik_{m}}-e^{-ik_{m^{\prime}}})\tilde{x}_{m}\tilde{x}_{m^{\prime}}\right].$ (110)